OpenStudy (anonymous):

Use differentials to approximate the maximum error propagated when calculating the volume of a sphere, if radius is 5 +/- .01 meters.

OpenStudy (anonymous):

@freckles @DanJS

OpenStudy (irishboy123):

do the differential first

OpenStudy (anonymous):

@IrishBoy123 differential of what?

OpenStudy (anonymous):

V=4/3πr3 ?

OpenStudy (anonymous):

@Nnesha

OpenStudy (gabebae):

What is the radius?

OpenStudy (anonymous):

5 + or - .01

OpenStudy (gabebae):

V=4/3πr^3 π or pi is equal to 3.14 22/7 and since you said the radius or 5 or -.01 you substitute that for r and cube it :)

OpenStudy (anonymous):

no radius is 5 +/- .01 so 2 radius here 5.01 and 4.99

OpenStudy (gabebae):

What's the full question?

OpenStudy (anonymous):

Use differentials to approximate the maximum error propagated when calculating the volume of a sphere, if radius is 5 +/- .01 meters. i typed the full question earlier. that is the full question

OpenStudy (gabebae):

This is calculus isn't it?

OpenStudy (anonymous):

yeah

OpenStudy (gabebae):

My friend is taking calculus but he isn't on at the moment. He lives in Autralia

OpenStudy (anonymous):

@DanJS is great but idk when he will be online.

OpenStudy (anonymous):

@Irishboy123

OpenStudy (irishboy123):

the volume is \(\large V = \dfrac{4}{3} \pi r^3\) the differential of that is what you need

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (anonymous):

the differential of dV=πx2dy? @Irishboy123?

OpenStudy (irishboy123):

you mentioned a sphere at the top of the thread \[\large V = \dfrac{4}{3} \pi r^3\] \[dV = ???\]

OpenStudy (anonymous):

4(pi)r^2

OpenStudy (anonymous):

@irishboy123?

OpenStudy (irishboy123):

what is \(\dfrac{d}{dr}( r^3)\) ??

OpenStudy (anonymous):

so 2 radius here 5.01 and 4.99